Problem stresstest SNF 2007-08-27 15:57:41: leaving in yellow until Uli checks it out

rissman at snf.stanford.edu rissman at snf.stanford.edu
Mon Aug 27 15:57:44 PDT 2007


Mary purchased a new unit.  With Eric's help I replaced the old unit (stage/laser head, motor controller) and confirmed operation.  A couple of comments:
1. The laser power (coming from the Coral box?) connector is a slightly different type and I had to remove the locking ears to get it to try to go in to the socket.  Even so it doesn't go in smoothly.  We may want to replace with a different (correct) connector.
2. Mary had hoped that the system would now have two laser frequencies (as the software indicates).  However, it only seems to have the 750 nm, not the 810 nm we were using (or viceversa, I get confused).
3. We have temporarily moved the pc on the table since the 64 pin cable is short.  We will try to get an extender.
System in yellow until Uli completely checks it out.




More information about the stresstest-pcs mailing list